If Mom was born in 1943 and had a son in 1978, in the year 2013 she would be twice his age. What is the formula to determine how old they would be when that would happen the next time? Thank you.
Found 2 solutions by Alan3354, Menjax:
Answer by Alan3354(69443)

If Mom was born in 1943 and had a son in 1978, in the year 2013 she would be twice his age. What is the formula to determine how old they would be when that would happen the next time?
------------
It won't happen again.
No next time.

Answer by Menjax(62)

m=s+35
m+35=s+70
m=2s
s=35 m=70
I think the question is asking when will the mom be double her sons age next?
s=40 m=80
formula is M=2s
y=ax+c
m=as+c
m=2s+c
80=80+0
m=2s